Based on the thermodynamic properties provided for water, determine the amount of energy needed for 2.60 kg of water to go from -18.0 °C to 77.0 °C.
(Q21)
Answer #1

The specific heat of water is 4.18 J/g.oC.

We have the relation

Q = m x Cp x dT

Here, m = 2.6 kg = 2600 g, Cp = 4.18 J/g.oC, dT = (77+18) = 95 oC

Q = 2600 x 4.18 x 95 = 1032.46 kJ

Amount of energy required = 1032.46 kJ